# Unifying the Perspectives of NLP and Software Engineering: A Survey on Language Models for Code

### v5

@@ Abstract
In this work we systematically review the recent advancements in code processing with language models, covering 50+ models, 30+ evaluation tasks, 170+ datasets, and 800 related works. We break down code processing models into general language models represented by the GPT family and specialized models that are specifically pretrained on code, often with tailored objectives. We discuss the relations and differences between these models, and highlight the historical transition of code modeling from statistical models and RNNs to pretrained Transformers and LLMs, which is exactly the same course that had been taken by NLP. We also discuss code-specific features such as AST, CFG, and unit tests, along with their application in training code language models, and identify key challenges and potential future directions in this domain.
@@ 1 Introduction
The application of pretrained Transformers in code processing, however, can be traced back to dates before decoder-only autoregressive models became dominant (Feng et al., 2020; Liu et al., 2020), and this domain is yet to witness a comprehensive review. In an attempt to bridge the gap between natural language processing (NLP) community and software engineering (SE) community on the topic of language model applications, we undertake a panoramic survey of language models for code in this work, covering 50+ models, 30+ downstream tasks, 170+ datasets, and 800 related works. We break down different categories of code language models, ranging from colossal models trained on general domains to tiny models trained specifically for code understanding or generation. We emphasize on the relations and differences between such models, and highlight the integration of code-specific features, such as abstract syntax trees or data flows, into language models, as well as the latest techniques adapted from NLP.
@@
SynCoBERT (Wang et al., 2021c), Code-MVP (Wang et al., 2022d), SCodeR (Li et al., 2022e),
CodeSage (Zhang et al., 2024),
@@
[CompCoder (Wang et al., 2022c), CodeRL (Le et al., 2022), PPOCoder (Shojaee et al., 2023), RLTF (Liu et al., 2023d),
@@
Thus, building on these works, we endeavor to unite the perspectives from both communities, and accentuate the integration between NLP and SE throughout the work. We make the key observation that advanced topics from language modeling have been recently introduced into code processing, including instruction tuning (Honovich et al., 2023; Xu et al., 2024; Luo et al., 2023), infilling objectives (Tay et al., 2023b; Li et al., 2023h; Rozière et al., 2023), recontemplation of scaling laws (Hoffmann et al., 2022; Gunasekar et al., 2023; Li et al., 2023i), architectural improvements (Shazeer, 2019; Su et al., 2024; Dao et al., 2022), and autonomous agents (Qian et al., 2023; Hong et al., 2023), while in return SE requirements are providing real-world testbeds for these technologies and driving the development of LLMs forward into production. We believe a systematic review of these advancements would benefit both communities.
@@
The rest of this work is organized following the taxonomy presented in Figure 1. In Section 2 we first contextualize the downstream tasks in code processing, highlighting the historical transition from various code understanding tasks to more practical text-to-code generation tasks. Then, in Section 3 we provide the preliminaries of language modeling and Transformer models, and in Section 4 we discuss the plethora of LLMs that have demonstrated coding ability. In Section 5 we review the specialized and often smaller models by their architecture, with special attention on the recent application of infilling objectives, instruction tuning, reinforcement learning, and engineering improvements. Then, in Section 6, we discuss unique features of code that are not available to natural languages but have been utilized to aid code processing. In Section 7, we review the most recent integration between LLMs and software development, before finally concluding this work in Section 8 and highlighting the current challenges in code processing.
2 Downstream Tasks in Code Processing
@@
Over the past decade, the software engineering community has found applications of code models in various downstream tasks. CodeXGLUE (Lu et al., 2021) consolidates many of such tasks into a single benchmark covering code understanding tasks such as clone detection, defect detection, and sequence-to-sequence generation tasks such as code repair, code translation, program synthesis, and code summarization. After Chen et al. (2021b) introduced HumanEval and Codex, however, text-to-code synthesis was brought into the spotlight in the NLP community and has since become a standard task for evaluating LLMs (Figure 2). Accordingly, we first briefly introduce each of the traditional tasks and the application of pretrained language models in them in Section 2.1, and provide a comprehensive list of related works for each task. Then, we review the evaluation metrics in Section 2.2 and investigate program synthesis in more detail in Section 2.3. Lastly, we also discuss the latest trend of repository-level coding in Section 2.4. In Appendix A, we list benchmarks for each downstream task.
@@
where level=3text width=6em,font=,
where level=4text width=8em,font=,
where level=5text width=13em,font=,
@@
[Code Evaluation, fill=black, fill opacity=0.7, text=white
[Text-to-Code, fill=yellow, fill opacity=0.5
[Code
Retrieval, fill=yellow, fill opacity=0.5
@@
RACS (Li et al., 2016), CodeMatcher (Liu et al., 2022a), fill=yellow, fill opacity=0.5
@@
TranCS (Sun et al., 2022), fill=yellow, fill opacity=0.5
@@
CDCS (Chai et al., 2022), CodeRetriever (Li et al., 2022d),
@@
CCT-LM (Sorokin et al., 2023), fill=yellow, fill opacity=0.5
[Text-to-SQL, fill=yellow, fill opacity=0.5
[, fill=yellow, fill opacity=0.5
@@
GNN (Bogin et al., 2019), TREQS (Wang et al., 2020b), fill=yellow, fill opacity=0.5
@@
Self-Debugging (Chen et al., 2023d), DIN-SQL (Pourreza & Rafiei, 2023),
@@
Guo et al. (2023a), DAIL-SQL (Gao et al., 2023a),
SQL-Encoder (Pourreza et al., 2024), fill=yellow, fill opacity=0.5
[Code
Synthesis, fill=yellow, fill opacity=0.5
Euphony (Lee et al., 2018), Neo (Feng et al., 2018), fill=yellow, fill opacity=0.5
LPN (Ling et al., 2016), NSPS (Parisotto et al., 2017),
DeepCoder (Balog et al., 2017), RobustFill (Devlin et al., 2017a),
(Yin & Neubig, 2017), ASN (Rabinovich et al., 2017),
NGDS (Kalyan et al., 2018), Bunel et al. (2018),
ReCode (Hayati et al., 2018), AutoPandas (Bavishi et al., 2019),
Wei et al. (2019), PlotCoder (Chen et al., 2021c),
Huang et al. (2022a), fill=yellow, fill opacity=0.5
TreeGen (Sun et al., 2020), REDCODER (Parvez et al., 2021),
Jigsaw (Jain et al., 2022), JuPyT5 (Chandel et al., 2022),
CodeT (Chen et al., 2023a), TiCoder (Lahiri et al., 2022),
AceCoder (Li et al., 2023e), Self-Debugging (Chen et al., 2023d),
ClarifyGPT (Mu et al., 2023),
MANGO (Chen et al., 2024), fill=yellow, fill opacity=0.5
[Math Programming, fill=yellow, fill opacity=0.5
[, fill=yellow, fill opacity=0.5
[, fill=yellow, fill opacity=0.5
MathQA-Python (Austin et al., 2021), Drori & Verma (2021),
Drori et al. (2022),
PAL (Gao et al., 2023b), PoT (Chen et al., 2023c),
He-Yueya et al. (2023),
GSM8K-Python (Chowdhery et al., 2023; Wang et al., 2023f), fill=yellow, fill opacity=0.5
[Code-to-Code, fill=cyan, fill opacity=0.4]
[Code-to-Text, fill=Orange, fill opacity=0.2]
[Code-to-Pattern, fill=VioletRed, fill opacity=0.2]
[Text-to-Text, fill=Dandelion, fill opacity=0.4]
2.1 Downstream Applications of Code Modeling
Following the custom in software engineering, we categorize the applications of code models according to their input/output modality, and break down these tasks into five families: text-to-code, code-to-code, code-to-text, code-to-pattern, and text-to-text. We note that this taxonomy is interleaved with the understanding-generation dichotomy in NLP, since each category may contain both understanding and generation tasks, as discussed in Section 2.1.6.
2.1.1 Text-to-Code
Text-to-code tasks take text as input, and output code. Related works are listed in Figure 3
@@ 2.1.1 Text-to-Code
- Code retrieval aims to retrieve relevant code given natural language queries, or to mine parallel text-code pairs from an unannotated corpus. This task is usually performed by computing a similarity metric between the embedding of query and candidate code, and the contextual embeddings produced by bidirectional language models - such as BERT - has proven to be extremely helpful. Grazia & Pradel (2023) and Xie et al. (2023a) provide comprehensive reviews on this topic.
@@
- Code synthesis aims to generate code (usually a function or a method) given a natural language description. This task can be viewed as an updated version of code retrieval using generative models instead of retrieval models. Statistical machine translation (SMT) and neural machine translation (NMT) models have been widely adopted for this task, often with enhanced decoders that leverage the unique grammatical rules of programming languages (Yin & Neubig, 2017; Rabinovich et al., 2017). Pretrained language models based on Transformer architecture, however, changed the game by directly generating the source code in the autoregressive language modeling style, even without task-specific finetuning (Chen et al., 2021b). We discuss this task in more detail in Section 2.3.
@@
- Text-to-SQL is a special (and arguably easier) case of code synthesis, where the model is tasked to generate SQL commands from natural language queries. It has been a topic of special interest due to SQL’s structured nature (when compared with general-purpose languages such as Python and C) and wide application in data management. We refer to Kumar et al. (2022); Deng et al. (2022a); Qin et al. (2022a); Katsogiannis-Meimarakis & Koutrika (2023) for surveys on this topic.
- Math programming is also a special case of code synthesis, where a language model is required to solve mathematical reasoning problems via generating code that will be executed by external interpreters. This task abstracts the reasoning process from numerical calculations, and is thus of special interest in evaluating LLMs.
2.1.2 Code-to-Code
@@
where level=3text width=8em,font=,
where level=4text width=8em,font=,
where level=5text width=11em,font=,
@@
[Code Evaluation, fill=black, fill opacity=0.7, text=white
[Text-to-Code, fill=yellow, fill opacity=0.5
[Code-to-Code, fill=cyan, fill opacity=0.2
[Code Search, fill=cyan, fill opacity=0.2
FaCoY (Kim et al., 2018), Aroma (Luan et al., 2019),
COSAL (Mathew & Stolee, 2021), fill=cyan, fill opacity=0.2
[, fill=cyan, fill opacity=0.2
Corder (Bui et al., 2021b), Cosco (Saieva et al., 2023), fill=cyan, fill opacity=0.2
@@
Completion, fill=cyan, fill opacity=0.2
Deep3 (Raychev et al., 2016a), Hellendoorn & Devanbu (2017), fill=cyan, fill opacity=0.2
Raychev et al. (2014), White et al. (2015),
Pointer Mixture Network (Li et al., 2018a), Parvez et al. (2018),
SLM (Alon et al., 2020), Pythia (Svyatkovskiy et al., 2019), NLM (Karampatsis et al., 2020), fill=cyan, fill opacity=0.2
IntelliCode (Svyatkovskiy et al., 2020), CugLM (Liu et al., 2020),
eWASH (Clement et al., 2021),
LongCoder (Guo et al., 2023b), fill=cyan, fill opacity=0.2
[Code
Translation, fill=cyan, fill opacity=0.2
@@
mppSMT (Nguyen et al., 2015), fill=cyan, fill opacity=0.2
Tree2Tree (Chen et al., 2018), Grammar Tree2Tree (Drissi et al., 2018), fill=cyan, fill opacity=0.2
@@
Self-Debugging (Chen et al., 2023d),
@@
Macedo et al. (2024), fill=cyan, fill opacity=0.2
[API
Mining, fill=cyan, fill opacity=0.2
RASH (Zhang et al., 2018), RAPIM (Alrubaye et al., 2020), fill=cyan, fill opacity=0.2
Nguyen et al. (2016), DeepAM (Gu et al., 2017),
JV2CS (Phan et al., 2017), Lu et al. (2017),
TL-CodeSum (Hu et al., 2018b),
BIKER (Huang et al., 2018), Chen et al. (2021a),
SAR (Bui et al., 2019), fill=cyan, fill opacity=0.2
HaPiM (Zhou et al., 2023b), fill=cyan, fill opacity=0.2
[Cloze Test, fill=cyan, fill opacity=0.2
[, fill=cyan, fill opacity=0.2
[, fill=cyan, fill opacity=0.2
CodeBERT (Feng et al., 2020),
Lu et al. (2021),
Puri et al. (2021),
JavaBERT (de Sousa & Hasselbring, 2021), fill=cyan, fill opacity=0.2
@@
Repair, fill=cyan, fill opacity=0.2
@@
Refactory (Hu et al., 2019), PyTER (Oh & Oh, 2022), fill=cyan, fill opacity=0.2
@@
CoCoNuT (Lutellier et al., 2020), DLFix (Li et al., 2020c),
@@
DEAR (Li et al., 2022f), fill=cyan, fill opacity=0.2

### v6

@@ Abstract
In this work we systematically review the recent advancements in software engineering with language models, covering 70+ models, 40+ evaluation tasks, 180+ datasets, and 900 related works. We break down code processing models into general language models represented by the GPT family and specialized models that are specifically pretrained on code, often with tailored objectives. We discuss the relations and differences between these models, and highlight the historical transition of code modeling from statistical models and RNNs to pretrained Transformers and LLMs, which is exactly the same course that had been taken by NLP. We also go beyond programming and review LLMs’ application in other software engineering activities including requirement engineering, testing, deployment, and operations in an endeavor to provide a global view of NLP in SE. We identify key challenges and potential future directions in this domain, and keep the survey open and updated on GitHub at https://github.com/codefuse-ai/Awesome-Code-LLM.
@@ 1 Introduction
The application of pretrained Transformers in software engineering, however, can be traced back to dates before decoder-only autoregressive models became dominant (Feng et al., 2020; Liu et al., 2020), and also goes beyond code processing (Arora et al., 2023; Feng et al., 2024; Ma et al., 2024). However, this domain is yet to witness a comprehensive review. In an attempt to bridge the gap between natural language processing (NLP) community and software engineering (SE) community on the topic of language model applications, we undertake a panoramic survey of language models for SE in this work, covering 70+ models, 40+ downstream tasks, 180+ datasets, and 900 related works. We break down different categories of code language models, ranging from colossal models trained on general domains to tiny models trained specifically for code understanding or generation. We emphasize on the relations and differences between such models, and highlight the integration of code-specific features, such as abstract syntax trees or data flows, into language models, as well as the latest techniques adapted from NLP.
@@
SynCoBERT (Wang et al., 2021d), Code-MVP (Wang et al., 2022e), SCodeR (Li et al., 2022f),
CodeSage (Zhang et al., 2024a),
@@
[CompCoder (Wang et al., 2022d), CodeRL (Le et al., 2022), PPOCoder (Shojaee et al., 2023), RLTF (Liu et al., 2023d),
@@
Thus, building on these works, we endeavor to unite the perspectives from both communities, and accentuate the integration between NLP and SE throughout the work. We make the key observation that advanced topics from language modeling have been recently introduced into code processing, including instruction tuning (Honovich et al., 2023; Xu et al., 2024; Luo et al., 2023), infilling objectives (Tay et al., 2023b; Li et al., 2023h; Rozière et al., 2023), recontemplation of scaling laws (Hoffmann et al., 2022; Gunasekar et al., 2023; Li et al., 2023i), architectural improvements (Shazeer, 2019; Su et al., 2024; Dao et al., 2022), and autonomous agents (Qian et al., 2023; Hong et al., 2023), while in return SE requirements are providing real-world testbeds for these technologies and driving the development of LLMs forward into production. We believe a systematic review of these advancements would benefit both communities. Furthermore, all of the existing reviews focus on programming-related tasks, while we are the first to explicitly go beyond programming and provide a global view of LLM applications in the full life cycle of software development, covering distinct stages such as requirement engineering, deployment, and operations.
@@
The rest of this work is organized following the taxonomy presented in Figure 1. In Section 2 we first contextualize the downstream tasks in software engineering, highlighting the focus on programming related tasks. Then, in Section 3 we provide the preliminaries of language modeling and Transformer models, and in Section 4 we discuss the plethora of LLMs that have demonstrated coding ability. In Section 5 we review the specialized and often smaller models by their architecture, with special attention on the recent application of infilling objectives, instruction tuning, reinforcement learning, and engineering improvements. Then, in Section 6, we discuss unique features of code that are not available to natural languages but have been utilized to aid code processing. In Section 7, we review the most recent integration between LLMs and software development, before finally concluding this work in Section 8 and highlighting the current challenges in code processing.
2 Downstream Tasks in Software Engineering
Over the past decade, the SE community has found applications of language models in various downstream tasks. CodeXGLUE (Lu et al., 2021) consolidates most of the code-related tasks into a single benchmark, while HumanEval Chen et al. (2021b) brought NL-to-code synthesis into the spotlight in the NLP community, which has since become a standard task for evaluating LLMs (Figure 2). However, other tasks, especially those not directly related to coding, have remained understudied.
@@ 2 Downstream Tasks in Software Engineering
In this section, we first briefly introduce each of the traditional SE tasks and the application of pretrained language models in them in 2.1, and provide a comprehensive list of related works for each task. Then, we review the evaluation metrics in 2.2 and investigate program synthesis in more detail in 2.3. Lastly, we also discuss the latest trend of repository-level coding in 2.4. In Appendix A, we list benchmarks for each downstream task.
2.1 Downstream Tasks in SE
The custom in software engineering is to categorize downstream tasks according to their input/output modality, such as NL-to-PL (also referred to as text-to-code) tasks and PL-to-NL (i.e. code-to-text) tasks (Lu et al., 2021). However, as we are dedicated to the full life cycle of software development, we adopt a different taxonomy in this work, and classify downstream tasks according to the stages in software development: requirement engineering (2.1.1), development (2.1.2), testing and analysis (2.1.3), deployment and operations (2.1.4), and lastly several novel tasks recently proposed for evaluating LLMs (2.1.5). We note that this taxonomy is interleaved with the understanding-generation dichotomy in NLP, since each category may contain both understanding and generation tasks, as discussed in 2.1.6.
2.1.1 Requirement Engineering
Requirement engineering refers to the specification, analysis, and validation of software requirements, software modeling, and UI/UX design. Related works are listed in Figure 3.
@@ 2.1.1 Requirement Engineering
where level=3text width=10.5em,font=,
where level=4text width=5.7em,font=,
where level=5text width=10.8em,font=,
@@
[NLP in SE, fill=black, fill opacity=0.7, text=white
[Requirement, fill=yellow, fill opacity=0.5
[Requirement Analysis, fill=yellow, fill opacity=0.5
ARSENAL (Ghosh et al., 2016),
CAR (Ferrari et al., 2014),
RETA (Arora et al., 2015a),
NARCIA (Arora et al., 2015b),
FENL (Bakar et al., 2016),
SNACC (Zhang & El-Gohary, 2017),
REGICE (Arora et al., 2017),
AUR-BoW (Lu & Liang, 2017),
ALERTme (Guzman et al., 2017),
SAFE (Johann et al., 2017),
OCLgen (Wang et al., 2018b),
SEMIOS (Mezghani et al., 2018),
Tiwari et al. (2019)
, fill=yellow, fill opacity=0.5
, fill=yellow, fill opacity=0.5
Sainani et al. (2020),
RE-BERT (de Araújo & Marcacini, 2021),
White et al. (2023),
Rodriguez et al. (2023),
Jain et al. (2023),
Arora et al. (2023),
Ronanki et al. (2023),
COREQQA (Abualhaija et al., 2023),
CompliAT (Arora et al., 2024)
, fill=yellow, fill opacity=0.5
[UI/UX
Design, fill=yellow, fill opacity=0.5
, fill=yellow, fill opacity=0.5
pix2code (Beltramelli, 2018),
Aşıroğlu et al. (2019),
sketch2code (Robinson, 2019)
, fill=yellow, fill opacity=0.5
PixelHelp (Li et al., 2020c),
HTLM (Aghajanyan et al., 2022),
Screen2Words (Wang et al., 2021b),
MarkupLM (Li et al., 2022b),
HTMLBERT (Xie et al., 2021a),
DOM-LM (Deng et al., 2022b),
WebFormer (Wang et al., 2022b),
PromptMaker (Jiang et al., 2022),
Wang et al. (2023a),
Gur et al. (2023),
MenuCraft (Kargaran et al., 2023),
Wang et al. (2023d),
GPTDroid (Liu et al., 2023m),
ViCT (Soselia et al., 2023),
UI grammar (Lu et al., 2023c),
Canvil (Feng et al., 2024)
, fill=yellow, fill opacity=0.5
[Model
Generation, fill=yellow, fill opacity=0.5
SUGAR (Kumar & Sanyal, 2008),
UMGAR (Deeptimahanti & Babar, 2009),
RACE (Ibrahim & Ahmad, 2010),
Herchi & Abdessalem (2012),
TRAM (Letsholo et al., 2013),
Visual Narrator (Robeer et al., 2016),
Arora et al. (2016),
Elallaoui et al. (2018),
Günes & Aydemir (2020),
Abdelnabi et al. (2020),
READ (Bashir et al., 2021),
Abdelnabi et al. (2021b)
, fill=yellow, fill opacity=0.5
, fill=yellow, fill opacity=0.5
Weyssow et al. (2022),
Chaaben et al. (2023),
Cámara et al. (2023),
Chen et al. (2023b),
Chen et al. (2023c),
Ferrari et al. (2024)
, fill=yellow, fill opacity=0.5
[Development, fill=Orange, fill opacity=0.2]
[Testing &
Analysis, fill=cyan, fill opacity=0.2]
[Deployment & Operations, fill=VioletRed, fill opacity=0.2]
[LLM Evaluation, fill=Dandelion, fill opacity=0.4]
- Requirement analysis refers to the process of extracting, summarizing, classifying, and validating software requirements. Most early works in this field rely on statistical NLP technologies such as POS (Part-of-Speech) tagging and dependency parsing, while more recent works utilize LLMs to classify, summarize, or extract requirement clauses. Zhao et al. (2020) provide a fine-grained survey on requirement analysis.
- UI/UX Design, short for User Interface and User Experience design, is a fundamental step in software development. While many works on UI design utilize computer vision technologies for layout design (Cheng et al., 2023; Kulkarni et al., 2023; Feng et al., 2023b), we only focus on the intersection of this task and NLP technologies, such as using language models to generate markup languages.
- Model generation aims to generate software models in modeling languages such as UML (Unified Modeling Language). Abdelnabi et al. (2021a) and Ahmed et al. (2022) provide comprehensive reviews on this task before the rise of LLMs.
{forest}
for tree=
forked edges,
draw,
rounded corners,
grow=east,
anchor=base west,
anchor=center,
reversed=true,
l sep=0.2cm,
where level=0font=,
where level=1text width=4.9em,font=,
where level=2text width=3.5em,font=,
where level=3text width=6.0em,font=,
where level=4text width=9.0em,font=,
where level=5text width=12.0em,font=,
for tree=
if level=0l sep+=0.15cm,
if level=1l sep+=0.15cm,
if level=2l sep+=0.03cm,
[NLP in SE, fill=black, fill opacity=0.7, text=white
[Requirement, fill=yellow, fill opacity=0.5]
[Development:
Code Search, fill=Orange, fill opacity=0.2
[NL-to-Code
Search, fill=Orange, fill opacity=0.2
@@
RACS (Li et al., 2016), CodeMatcher (Liu et al., 2022a), fill=Orange, fill opacity=0.2
@@
TranCS (Sun et al., 2022), fill=Orange, fill opacity=0.2
@@
CDCS (Chai et al., 2022), CodeRetriever (Li et al., 2022e),
@@
CCT-LM (Sorokin et al., 2023), fill=Orange, fill opacity=0.2
[Code-to-Code
Search, fill=Orange, fill opacity=0.2
FaCoY (Kim et al., 2018), Aroma (Luan et al., 2019),
COSAL (Mathew & Stolee, 2021), fill=Orange, fill opacity=0.2
[, fill=Orange, fill opacity=0.2
Corder (Bui et al., 2021b), Cosco (Saieva et al., 2023), fill=Orange, fill opacity=0.2
[API
Mining, fill=Orange, fill opacity=0.2
RASH (Zhang et al., 2018), RAPIM (Alrubaye et al., 2020), fill=Orange, fill opacity=0.2
Nguyen et al. (2016), DeepAM (Gu et al., 2017),
JV2CS (Phan et al., 2017), Lu et al. (2017),
TL-CodeSum (Hu et al., 2018b),
BIKER (Huang et al., 2018), Chen et al. (2021a),
SAR (Bui et al., 2019), fill=Orange, fill opacity=0.2
HaPiM (Zhou et al., 2023b),
Code2API (Mai et al., 2024)
, fill=Orange, fill opacity=0.2
[Development:
Code Generation, fill=Orange, fill opacity=0.2
[Program
Synthesis, fill=Orange, fill opacity=0.2
Euphony (Lee et al., 2018), Neo (Feng et al., 2018), fill=Orange, fill opacity=0.2
LPN (Ling et al., 2016), NSPS (Parisotto et al., 2017),
DeepCoder (Balog et al., 2017), RobustFill (Devlin et al., 2017a),
(Yin & Neubig, 2017), ASN (Rabinovich et al., 2017),
NGDS (Kalyan et al., 2018), Bunel et al. (2018),
ReCode (Hayati et al., 2018), AutoPandas (Bavishi et al., 2019),
Wei et al. (2019), PlotCoder (Chen et al., 2021c),
Huang et al. (2022a), fill=Orange, fill opacity=0.2
TreeGen (Sun et al., 2020), REDCODER (Parvez et al., 2021),
Jigsaw (Jain et al., 2022), JuPyT5 (Chandel et al., 2022),
CodeT (Chen et al., 2023a), TiCoder (Lahiri et al., 2022),
AceCoder (Li et al., 2023e), Self-Debugging (Chen et al., 2023f),
ClarifyGPT (Mu et al., 2023b),
MANGO (Chen et al., 2024), fill=Orange, fill opacity=0.2
[Code
Completion, fill=Orange, fill opacity=0.2
Deep3 (Raychev et al., 2016a), Hellendoorn & Devanbu (2017), fill=Orange, fill opacity=0.2
Raychev et al. (2014), White et al. (2015),
Pointer Mixture Network (Li et al., 2018a), Parvez et al. (2018),
SLM (Alon et al., 2020), Pythia (Svyatkovskiy et al., 2019), NLM (Karampatsis et al., 2020), fill=Orange, fill opacity=0.2
IntelliCode (Svyatkovskiy et al., 2020), CugLM (Liu et al., 2020),
eWASH (Clement et al., 2021),
LongCoder (Guo et al., 2023b), fill=Orange, fill opacity=0.2
[Code
Infilling, fill=Orange, fill opacity=0.2
[, fill=Orange, fill opacity=0.2
[, fill=Orange, fill opacity=0.2
StarCoder (Li et al., 2023h), Code LLaMA (Rozière et al., 2023), fill=Orange, fill opacity=0.2
[Text-to-SQL, fill=Orange, fill opacity=0.2
[, fill=Orange, fill opacity=0.2, text width=3em
@@
GNN (Bogin et al., 2019), TREQS (Wang et al., 2020b), fill=Orange, fill opacity=0.2
@@
DIN-SQL (Pourreza & Rafiei, 2023),
@@
Guo et al. (2023a), DAIL-SQL (Gao et al., 2023a)
, fill=Orange, fill opacity=0.2, text width=15em
[Development:
Code Editing, fill=Orange, fill opacity=0.2]
[Development:
Code Suggestion, fill=Orange, fill opacity=0.2]
[Development:
Code Explanation, fill=Orange, fill opacity=0.2]
[Testing &
Analysis, fill=cyan, fill opacity=0.2]
[Deployment & Operations, fill=VioletRed, fill opacity=0.2]
[LLM Evaluation, fill=Dandelion, fill opacity=0.4]
2.1.2 Development
The development phase is the stage where most programming activities take place, and also the stage where LLM applications are most abundant. We further categorize this stage into different activities such as code search, code generation, code editing, code suggestion, and code explanation. Related works are listed in Figure 4 and 5.
Code Search
@@ 2.1.2 Development
- NL-to-code search, also referred to as text-to-code search, aims to retrieve relevant code given natural language queries, or to mine parallel text-code pairs from an unannotated corpus. This task is usually performed by computing a similarity metric between the embedding of query and candidate code, and the contextual embeddings produced by bidirectional language models - such as BERT - has proven to be extremely helpful. Grazia & Pradel (2023) and Xie et al. (2023a) provide comprehensive reviews on this topic.
- Code-to-code search is a similar task where the input is an existing code snippet, often in a different programming language from the target. Code-to-code search can be reformulated as finding clones of the query in the pool of targets, and is thus equivalent to clone detection to some extent.
- API mining refers to the process of finding similar APIs in different libraries, potentially in different programming languages. API mining is traditionally tackled by computing similarity metrics between source and target APIs using information retrieval models, but as generative models become ever more capable, it is also worth exploring to directly generate the target API as a sequence-to-sequence task. Another closely related task is idiom mining (Allamanis & Sutton, 2014), where the objective is to discover commonly used code patterns, which exposes the potential need for new APIs (Sivaraman et al., 2022).
Code Generation
@@
- Program synthesis aims to generate code (usually a function or a method) given a natural language description. This task can be viewed as an updated version of NL-to-code retrieval using generative models instead of retrieval models. Statistical machine translation (SMT) and neural machine translation (NMT) models have been widely adopted for this task, often with enhanced decoders that leverage the unique grammatical rules of programming languages (Yin & Neubig, 2017; Rabinovich et al., 2017). Pretrained language models based on Transformer architecture, however, changed the game by directly generating the source code in the autoregressive language modeling style, even without task-specific finetuning (Chen et al., 2021b). We discuss this task in more detail in Section 2.3.
- Code completion aims to complete a piece of code given its prefix, and remains to date one of the most popular applications of code language models in IDEs. This is essentially language modeling applied to code (which we dub “code modeling”), and related technologies have been progressively introduced: n-gram, RNN, and Transformer. However, due to the structured nature of programming languages, many early works found grammar-aided statistical models to perform better (Bielik et al., 2016; Hellendoorn & Devanbu, 2017), and neural models only became dominant after 2018.
- Code infilling is another recently proposed task, after fill-in-the-middle pretraining (Bavarian et al., 2022) became popular. It is a generalization of code completion, where not only the left context but also the right context is given.
@@
- Text-to-SQL is a special case of program synthesis, where the model is tasked to generate SQL commands from natural language queries. It has been a topic of special interest in the NLP community (as can be seen from Figure 4), and one hypothesis for this is that SQL’s declarative nature and restricted grammar make it easier to optimize compared with imperative programming languages such as C and Python (Marcus et al., 2019; 2020). We refer to Kumar et al. (2022); Deng et al. (2022a); Qin et al. (2022a); Katsogiannis-Meimarakis & Koutrika (2023) for surveys on this topic.
@@
where level=3text width=6.0em,font=,
where level=4text width=9.0em,font=,
where level=5text width=12.0em,font=,
@@
[NLP in SE, fill=black, fill opacity=0.7, text=white
[Requirement, fill=yellow, fill opacity=0.5]
[Development:
Code Search, fill=Orange, fill opacity=0.2]
[Development:
Code Editing, fill=Orange, fill opacity=0.2
@@
Translation, fill=Orange, fill opacity=0.2
@@
mppSMT (Nguyen et al., 2015), fill=Orange, fill opacity=0.2
Tree2Tree (Chen et al., 2018), Grammar Tree2Tree (Drissi et al., 2018), fill=Orange, fill opacity=0.2
@@
Self-Debugging (Chen et al., 2023f),
@@
Macedo et al. (2024), fill=Orange, fill opacity=0.2
@@
Repair, fill=Orange, fill opacity=0.2
@@
Refactory (Hu et al., 2019), PyTER (Oh & Oh, 2022), fill=Orange, fill opacity=0.2
@@
CoCoNuT (Lutellier et al., 2020), DLFix (Li et al., 2020d),
@@
DEAR (Li et al., 2022g), fill=Orange, fill opacity=0.2